OpenStudy (vera_ewing):

You are going to design an advertisement for a new polynomial identity that you are going to invent. Your goal for this activity is to demonstrate the proof of your polynomial identity through an algebraic proof and a numerical proof in an engaging way! Make it so the whole world wants to purchase your polynomial identity and can't imagine living without it!

OpenStudy (vera_ewing):

You may do this by making a flier, a newspaper or magazine advertisement, making an infomercial video or audio recording, or designing a visual presentation for investors through a flowchart or PowerPoint. You must: Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step Demonstrate that your polynomial identity works on numerical relationships

OpenStudy (vera_ewing):

@SolomonZelman

OpenStudy (vera_ewing):

@uri

OpenStudy (campbell_st):

have you come up with a new polynomial identity..?

OpenStudy (vera_ewing):

Yes

OpenStudy (vera_ewing):

It says: 1. Label and display your new polynomial identity 2. Prove that it is true through an algebraic proof, identifying each step 3. Demonstrate that your polynomial identity works on numerical relationships

OpenStudy (vera_ewing):

How about (6-2) Would that work?

OpenStudy (campbell_st):

well its the difference of 2 degree 1 polynomials...

OpenStudy (vera_ewing):

Ok so what would you do?

OpenStudy (campbell_st):

how would you prove it algebraically

OpenStudy (vera_ewing):

Wait so does mine work? We can use (6-2) ?

OpenStudy (campbell_st):

yes but you need to prove it....

OpenStudy (vera_ewing):

Ok so what would I say?

OpenStudy (campbell_st):

well you could say P(x) = a and then H(x) = b then the difference between the polynomials P(x) - H(x) = a - b start with that... how do you prove subtraction works...

OpenStudy (vera_ewing):

@campbell_st Ok so is that it?

OpenStudy (vera_ewing):

@SolomonZelman

OpenStudy (campbell_st):

well I'm not sure...

OpenStudy (vera_ewing):

oh...ok. I will ask someone else. Thanks for helping

OpenStudy (vera_ewing):

Do you know how to do this one?

OpenStudy (solomonzelman):

Well, I would try to pick something a little harder. (Rather than subtracting two constant polynomials.)

OpenStudy (vera_ewing):

Ok what would you do?

OpenStudy (solomonzelman):

I am not even sure what the instructions exactly are for this assignment, but I would (perhaps) try to show how to find x-intercepts for any quadratic polynomial \(\large\color{black}{ \displaystyle y=\color{red}{\rm a} x^2+\color{green}{\rm b}x+\color{blue}{\rm c} }\).

OpenStudy (vera_ewing):

1. Label and display your new polynomial identity 2. Prove that it is true through an algebraic proof, identifying each step 3. Demonstrate that your polynomial identity works on numerical relationships

OpenStudy (solomonzelman):

I mean I would like you to come up with the idea for the assignment though. (something from the section which this is coming from)

OpenStudy (vera_ewing):

(x-y) was an example of what I could use.

OpenStudy (solomonzelman):

(x-y)=constant (though)

OpenStudy (solomonzelman):

like x-y=4

OpenStudy (solomonzelman):

that is a polynomial, but what would your identity then be ?

OpenStudy (solomonzelman):

Are you allowed to prove rules that people have already come up with ?

OpenStudy (solomonzelman):

(that **mathematicians** have already come up with )